### v2

# Equivalence of $f(Q)$ cosmology with quintom-like scenario: the phantom field as effective realization of the non-trivial connection

Abstract
We show that $f(Q)$ cosmology with a non-trivial connection, namely the Connection II of the literature, is dynamically equivalent with a quintom-like model. In particular, we show that the scalar field arising from the non-linear $f(Q)$ form, and the scalar field associated to the non-trivial connection, are combined to provide one canonical and one phantom field in the minisuperspace Lagrangian. Hence, the combination of two well-behaved fields appears effectively as a phantom field. This is fundamentally different from the usual approach in phantom and quintom cosmology, in which the phantom field is introduced ad hoc, and thus it may open a new way to handle the phantom fields, namely as effective realizations of the connection-related canonical fields.
pacs:
04.50.Kd, 98.80.-k, 95.36.+x
I Introduction
In an effort to address certain theoretical issues of General Relativity, such as its non-renormalizability Stelle:1976gc , as well as the cosmological constant problem Weinberg:1988cp , but also to provide potential ways to alleviate various observational tensions Perivolaropoulos:2021jda ; Abdalla:2022yfr , a wide range of modified gravity theories have been developed CANTATA:2021asi . There exist various avenues for constructing such modifications. A common approach begins with the standard curvature-based formulation of gravity, leading to theories such as $f(R)$ gravity Starobinsky:1980te ; DeFelice:2010aj , $f(G)$ gravity Nojiri:2005jg , $f(P)$ gravity Erices:2019mkd , and Lovelock gravity Lovelock:1971yv , among others. Alternatively, one may consider the torsional formulation of gravity, which gives rise to frameworks like $f(\mathbb{T})$ gravity Bengochea:2008gz ; Cai:2015emx and $f(\mathbb{T},\mathbb{T}_{G})$ gravity Kofinas:2014owa . However, there is a third powerful path that has attracted the interest of the community, and involves the use of non-metricity as the fundamental geometric quantity Nester:1998mp ; BeltranJimenez:2017tkd , leading to the construction of $f(Q)$ gravity theories BeltranJimenez:2019tme ; Anagnostopoulos:2021ydo ; Lazkoz:2019sjl ; Lu:2019hra ; Mandal:2020buf ; Ayuso:2020dcu ; Frusciante:2021sio ; Ferreira:2022jcd ; Gadbail:2022jco ; Sarmah:2023oum ; Khyllep:2021pcu ; Barros:2020bgg ; De:2022jvo ; Solanki:2022ccf ; DAmbrosio:2021zpm ; Li:2021mdp ; Dimakis:2021gby ; Kar:2021juu ; Wang:2021zaz ; Quiros:2021eju ; Mandal:2021bpd ; Albuquerque:2022eac ; Anagnostopoulos:2022gej ; Arora:2022mlo ; Pati:2022dwl ; Dimakis:2022wkj ; DAgostino:2022tdk ; Narawade:2022cgb ; Emtsova:2022uij ; Bahamonde:2022cmz ; Sokoliuk:2023ccw ; Shaikh:2023tii ; Dimakis:2023uib ; Koussour:2023rly ; Najera:2023wcw ; Atayde:2023aoj ; Shabani:2023xfn ; Tayde:2023xbm ; De:2023xua ; Junior:2024xmm ; Millano:2024rog ; Yang:2024tkw ; Guzman:2024cwa ; Alwan:2024lng ; Shaily:2024tmx ; Moreira:2024unj ; Dubey:2024gxa ; Mushtaq:2025jjj ; Nashed:2025usa ; Papagiannopoulos:2025uix ; Moreira:2025rxp ; Dimakis:2025jrl ; youri (for a review see Heisenberg:2023lru ).
On the other hand, the cosmological constant problem, as well as the observational issues of the Standard Model of Cosmology, may be addressed maintaining general relativity as the underlying gravitational theory but introducing extra scalar fields, with the notable examples being the inflaton Lidsey:1995np and the quintessence Ratra:1987rm . Although the scalar fields are typically considered as canonical ones, the need to describe the “phantom” regime, namely to obtain equation-of-state parameters smaller than the cosmological constant value $-1$ led to the wide use of phantom fields too, namely fields that have a negative kinetic energy in the Lagrangian Singh:2003vx . Despite the fact that the phantom fields can have many interesting cosmological implications and can lead to viable phenomenology Dabrowski:2003jm ; Chimento:2003qy ; Sami:2003xv ; Gonzalez-Diaz:2003bwh ; Caldwell:2003vq ; Elizalde:2004mq ; Perivolaropoulos:2004yr ; Nojiri:2005pu ; Faraoni:2005gg ; Nojiri:2005sx ; Gannouji:2006jm ; Chen:2008ft ; Elizalde:2008yf ; Saridakis:2009pj ; Astashenok:2012tv , their introduction is ad hoc, and the fact that their energy is unbounded from below makes them unstable at the quantum level Schon:1981vd ; Witten:1981mf ; Cline:2003gs ; Saridakis:2008fy . However, although there have been efforts to construct a phantom theory consistent with quantum field theory, for instance trying to make the phantom fields to arise as effective descriptions of higher-dimensional canonical theories Nojiri:2003vn ; Nojiri:2003ag , the whole issue is still open. Finally, the simultaneous consideration of one canonical and one phantom field, gives rise to the two-field, quintom model, which shares the advantages of both quintessence and phantom cosmology and moreover is able to describe the phantom-divide crossing Cai:2009zp ; Feng:2004ad ; Guo:2004fq ; Zhao:2005vj ; Zhao:2006mp ; Cai:2006dm ; Guo:2006pc ; Cai:2007qw . We mention here that the phantom-divide crossing is strongly favoured by various datasets, among others by the recently released DESI DR2 data Lodha:2025qbg .
In this work we show that $f(Q)$ cosmology with a non-trivial connection, namely the Connection II of the literature, is dynamically equivalent with a quintom-like model. In particular, we show that the scalar field arising from the non-linear $f(Q)$ form, and the scalar field associated to the non-trivial connection, are combined to give one canonical and one phantom field in the minisuperspace Lagrangian. Hence, the combination of two well-behaved fields appears effectively as a phantom field. This may open a new way to handle the phantom fields, namely as effective realizations of the connection-related canonical fields.
The plan of the work is the following: In Section II we review the basics of $f(Q)$ gravity and cosmology, presenting also the minisuperspace approach. In Section III we present the quintom scenario, namely the simultaneous consideration of a canonical and a phantom scalar field. Then in Section IV we show the dynamical equivalence of the two scenarios. Finally, Section V is devoted to the Conclusions.
II $f(Q)$ gravity and cosmology
In this section we intrdocue the $f(Q)$ gravity and we apply it in a cosmological framework.
II.1 $f(Q)$ gravity
We start by presenting the basics of symmetric teleparallel geometry. We introduce a general affine connection as
$\Gamma_{\mu\nu}^{\alpha}=\hat{\Gamma}_{\mu\nu}^{\alpha}+K_{\,\,\mu\nu}^{\alpha}+L_{\,\,\mu\nu}^{\alpha},$ (1)
with $\hat{\Gamma}_{\mu\nu}^{\alpha}$ is the Levi-Civita connection, $K_{\,\,\mu\nu}^{\alpha}=\frac{1}{2}T_{\,\,\mu\nu}^{\alpha}+T_{(\mu\,\,\,\nu)}^{\,\,\,\alpha}$ the contortion tensor related to the torsion tensor $T_{\,\,\mu\nu}^{\alpha}$, and with $L_{\,\,\mu\nu}^{\alpha}=\frac{1}{2}Q_{\,\,\mu\nu}^{\alpha}-Q_{(\mu\,\,\,\nu)}^{\,\,\,\alpha}$ the disformation tensor which arises from the non-metricity tensor
$Q_{\alpha\mu\nu}\equiv\nabla_{\alpha}g_{\mu\nu},$ (2)
with $g_{\mu\nu}$ the metric (throughout the manuscript Greek indices denote coordinate space). In terms of the affine connection we can write the torsion, curvature and nonmetricity tensors as
$T^{\lambda}{}_{\mu\nu}\equiv\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\lambda}{}_{\nu\mu}\,$
$R^{\sigma}{}_{\rho\mu\nu}\equiv\partial_{\mu}\Gamma^{\sigma}{}_{\nu\rho}-\partial_{\nu}\Gamma^{\sigma}{}_{\mu\rho}+\Gamma^{\alpha}{}_{\nu\rho}\Gamma^{\sigma}{}_{\mu\alpha}-\Gamma^{\alpha}{}_{\mu\rho}\Gamma^{\sigma}{}_{\nu\alpha}$
$Q_{\rho\mu\nu}\equiv\nabla_{\rho}g_{\mu\nu}=\partial_{\rho}g_{\mu\nu}-\Gamma^{\beta}{}_{\rho\mu}g_{\beta\nu}-\Gamma^{\beta}{}_{\rho\nu}g_{\mu\beta}\,.$ (3)
In general relativity one uses Riemannian geometry, in which case the gravitational Lagrangian is just the Ricci scalar derived from contractions of the Riemann tensor. Similarly, in the Teleparallel Equivalent of General Relativity one uses the Weitzenböck geometry, in which case the gravitational Lagrangian is just the torsion scalar derived from contractions of the torsion tensor. Hence, one can obtain a third description of gravity based on nonmetricity and the framework of symmetric teleparallel geometry.
In particular, one can use as gravitational Lagrangian the non-metricity scalar arising from contractions of the non-metricity tensor, namely
$Q=-\frac{1}{4}Q_{\alpha\beta\gamma}Q^{\alpha\beta\gamma}+\frac{1}{2}Q_{\alpha\beta\gamma}Q^{\gamma\beta\alpha}+\frac{1}{4}Q_{\alpha}Q^{\alpha}-\frac{1}{2}Q_{\alpha}\tilde{Q}^{\alpha}\,,$ (4)
with $Q_{\alpha}\equiv Q_{\alpha\ \mu}^{\ :\mu}\,,$ and $\tilde{Q}^{\alpha}\equiv Q_{\mu}^{\ :\mu\alpha}\,.$ Thus, similarly to the $f(R)$ and $f(\mathbb{T})$ extensions of the corresponding theories, one can extend the Lagrangian of the symmetric teleparallel equivalent of general relativity to an arbitrary function, obtaining $f(Q)$ gravity, with action BeltranJimenez:2017tkd :
$S=\frac{1}{2\kappa}\int{\mathrm{d}}^{4}x\sqrt{-g}f(Q)+S_{m},$ (5)
with $\kappa=8\pi G$ the gravitational constant, and where we have also included the matter sector action. Note that the Symmetric Teleparallel Equivalent of General Relativity, with or without cosmological constant, is obtained for linear function $f$, that is $f_{,QQ}\left(Q\right)=0$, or when $Q=$const.
Finally, performing variation yields the field equations of $f(Q)$ gravity as BeltranJimenez:2019tme :
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{2}{\sqrt{-g}}\nabla_{\lambda}(\sqrt{-g}f_{Q}P^{\lambda}{}_{\mu\nu})-\frac{1}{2}fg_{\mu\nu}$
$\ \ \ +f_{Q}(P_{\nu\rho\sigma}Q_{\mu}{}^{\rho\sigma}-2P_{\rho\sigma\mu}Q^{\rho\sigma}{}_{\nu})=\kappa T_{\mu\nu}^{m},$ (6)
with $T_{\mu\nu}^{m}$ the matter energy-momentum tensor, $f_{Q}=\partial f/\partial Q$, and where the superpotential $P^{\lambda}{}_{\mu\nu}$ is written as
$P^{\lambda}{}_{\mu\nu}=\frac{1}{4}\left(-2L^{\lambda}{}_{\mu\nu}+Q^{\lambda}g_{\mu\nu}-\tilde{Q}^{\lambda}g_{\mu\nu}-\frac{1}{2}\delta_{\mu}^{\lambda}Q_{\nu}-\frac{1}{2}\delta_{\nu}^{\lambda}Q_{\mu}\right).$ (7)
Additionally, variation of the action with respect to the general affine connection leads the connection equation as
$\nabla_{\mu}\nabla_{\nu}(\sqrt{-g}f_{Q}P^{\nu\mu}{}_{\lambda})=0\,,$ (8)
where we have assumed that the matter Lagrangian does not depend on the connection.
II.2 $f(Q)$ cosmology
We proceed to the application of $f(Q)$ gravity at a cosmological framework. We impose the flat Friedmann-Robertson-Walker (FRW) metric
$ds^{2}=-N(t)^{2}dt^{2}+a\left(t\right)^{2}\left(dr^{2}+r^{2}\mathrm{d}\theta^{2}+r^{2}\sin^{2}\theta\mathrm{d}\phi^{2}\right),$
where $a(t)$ is the scale factor, and $N(t)$ is the lapse function. In this case, the requirement the connection to be symmetric, flat and to inherit the symmetries of the background spacetime, leads to three different families of connections Hohmann:2020zre ; Hohmann:2021ast ; DAmbrosio:2021pnd ; Paliathanasis:2023nkb ; Paliathanasis:2015aos ; Dimakis:2016mip ; Heisenberg:2022mbo ; Shi:2023kvu . In particular, the nonzero connection components for a flat FRW metric are
$\Gamma_{~{}tt}^{t}$ $=K_{1}\,,~{}~{}\Gamma_{~{}rr}^{t}=K_{2}\,,~{}~{}\Gamma_{~{}\theta\theta}^{t}=K_{2}r^{2}\,,$
$\Gamma_{~{}tr}^{r}$ $=\Gamma_{~{}rt}^{r}=\Gamma_{~{}t\theta}^{\theta}=\Gamma_{~{}\theta t}^{\theta}=\Gamma_{~{}t\phi}^{\phi}=\Gamma_{~{}\phi t}^{\phi}=K_{3}\,,$
$\Gamma_{~{}r\theta}^{\theta}$ $=\Gamma_{~{}\theta r}^{\theta}=\Gamma_{~{}r\phi}^{\phi}=\Gamma_{~{}\phi r}^{\phi}=\frac{1}{r}\,,~{}~{}\Gamma_{~{}\theta\theta}^{r}=-r\,,$
$~{}~{}\Gamma_{~{}\phi\phi}^{r}$ $=-r\sin^{2}\theta\,,~{}~{}\Gamma_{~{}\phi\phi}^{t}=K_{2}r^{2}\sin^{2}\theta\,,$
$\Gamma_{~{}\phi\theta}^{\phi}$ $=\Gamma_{~{}\theta\phi}^{\phi}=\cot\theta\,,~{}~{}\Gamma_{~{}\phi\phi}^{\theta}=-\sin\theta\cos\theta\,,$ (9)
where $K_{1}(t),K_{2}(t),K_{3}(t)$ are functions of time, classified as
Connection $\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{Connection I}\!:K_{1}=\gamma(t),K_{2}=0,K_{3}=0,$ (10)
Connection $\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{Connection II}\!:K_{1}=\frac{\dot{\gamma}(t)}{\gamma(t)}+\gamma(t),K_{2}=0,K_{3}=\gamma(t),$ (11)
Connection $\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{Connection III}\!:K_{1}=-\frac{\dot{\gamma}(t)}{\gamma(t)},K_{2}=\gamma(t),K_{3}=0,$ (12)
where $\gamma(t)$ a function of time. Note that in the case of non-flat FRW geometry there is another connection, namely Connection IV, which however recovers Connection III in the flat case Hohmann:2020zre ; Hohmann:2021ast ; DAmbrosio:2021pnd ; Paliathanasis:2023nkb ; Paliathanasis:2015aos ; Dimakis:2016mip ; Heisenberg:2022mbo ; Shi:2023kvu . Connection I has been widely studied in cosmological studies, however the resulting background gravitational field equations are equivalent to that of $f(\mathbb{T})$-teleparallel gravity as discussed in Paliathanasis:2023nkb .
In this work we consider the Connection II above. Thus, the nonmetricity scalar is defined as
$Q\left(\Gamma\right)=-6H^{2}+\frac{3\gamma}{N}\left(3H-\frac{\dot{N}}{N^{2}}\right)+\frac{3\dot{\gamma}}{N^{2}}.$ (13)
Additionally, the general field equations (6) give rise to the Friedmann equations as
$3H^{2}f^{\prime}(Q)+\frac{1}{2}\left(f(Q)-Qf^{\prime}(Q)\right)+\frac{3\gamma\dot{Q}f^{\prime\prime}(Q)}{2N^{2}}=\kappa\rho_{m},$ (14)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!-\frac{2}{N}\frac{d}{dt}\left(f^{\prime}(Q)H\right)-3H^{2}f^{\prime}(Q)-\frac{1}{2}\left(f(Q)-Qf^{\prime}(Q)\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{3\gamma\dot{Q}f^{\prime\prime}(Q)}{2N^{2}}=\kappa p_{m},$ (15)
where primes denote derivatives with respect to $Q$.
II.3 Minisuperspace description
Let us now formulate $f(Q)$ cosmology in the mini-superspace framework. For simplicity in the following we set the gravitational constant $\kappa=1$ and we neglect the matter content. Consider the action in $f\left(Q\right)$ gravity, where we introduce the Lagrange multiplier $\lambda,~{}$such that
$S_{f\left(Q\right)}=\int\sqrt{-g}d^{4}x\left[f\left(Q\right)-\lambda\left(Q-\mathcal{Q}\right)\right],$ (16)
where $Q\equiv\mathcal{Q}\left(N,\dot{N},a,\dot{a},\gamma,\gamma\right)$ is the expression of the nonmetricity scalar. Variation with respect to $Q$, gives $\lambda=f^{\prime}\left(Q\right)$.
We replace in the action integral, and for the connection choice (11) we discussed in the previous subsection, it follows
[…]
Integrating by parts the last term yields
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\int dt\left[\frac{3}{N}f^{\prime}\left(Q\right)a^{3}\dot{\gamma}\right]=-\int dt\Big{[}-\frac{3\dot{N}}{N^{2}}f^{\prime}\left(Q\right)a^{3}\gamma$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{9}{N}f^{\prime}\left(Q\right)a^{2}\dot{a}\gamma+\frac{3}{N}f^{\prime\prime}\left(Q\right)a^{3}\dot{Q}\gamma\Big{]}$
[…]
Therefore, the point-like Lagrangian of the field equations is
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!L\left(\Gamma_{2}\right)=-\frac{3a\dot{a}^{2}f^{\prime}(Q)}{N}+\frac{N}{2}a^{3}\left[f(Q)-Qf^{\prime}(Q)\right]$
$\ \ -\frac{3a^{3}\dot{\psi}\dot{Q}f^{\prime\prime}(Q)}{2N},$ (18)
with the use of $\gamma=\dot{\psi}$.
We proceed by defining $\varphi=f^{\prime}\left(Q\right)$. Hence, we end with the scalar-tensor Lagrangian
$L\left(\Gamma^{B}\right)=-\frac{3}{N}\varphi a\dot{a}^{2}-\frac{3}{2N}a^{3}\dot{\varphi}\dot{\psi}-Na^{3}V\left(\varphi\right),$ (19)
where $V(\varphi)$ is the effective scalar field potential, related to the $f\left(Q\right)$ function as
$V\left(f^{\prime}(Q)\right)=\frac{1}{2}\left[Qf^{\prime}(Q)-f(Q)\right].$ (20)
III Quintom scenario
In this section we will briefly review the quintom cosmology. Since we are interested in comparing with the minisuperspace description of $f(Q)$ cosmology, we also set the gravitational constant $\kappa=1$ and we neglect the matter content.
The quintom scenario is based on the simultaneous consideration of two scalar fields: $\phi$, which is canonical, and $\sigma$, which is phantom Feng:2004ad . In particular, the action is written as Cai:2009zp :
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=\int d^{4}x\sqrt{-g}\bigg{[}R-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$
$\ \ \ \ \ \ \ \,+\frac{1}{2}g^{\mu\nu}\partial_{\mu}\sigma\partial_{\nu}\sigma+V\left(\phi,\sigma\right)\bigg{]},$ (21)
with $V\left(\phi,\sigma\right)$ the potential (one can split it in two separate potentials, such as $V\left(\phi,\sigma\right)=V_{\phi}(\phi)+V_{\sigma}(\sigma)$, however this is not necessary.
In a flat FRW geometry, variation of the above action with respect to the metric gives the Friedmann equations Cai:2009zp :
$H^{2}=\frac{1}{3}\Big{[}\frac{1}{2N^{2}}\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)+V\left(\phi,\sigma\right)\Big{]},$ (22)
$\dot{H}=-\frac{1}{2}\Big{[}\frac{1}{N^{2}}\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)\Big{]}.$ (23)
Additionally, variation of the action with respect to the scalar fields gives the Klein-Gordon equations as
$\frac{1}{N}\left(\frac{\dot{\phi}}{N}\right)^{\cdot}+\frac{3}{N}H\dot{\phi}+\frac{\partial V\left(\phi,\sigma\right)}{\partial\phi}=0$ (24)
$\frac{1}{N}\left(\frac{\dot{\sigma}}{N}\right)^{\cdot}+\frac{3}{N}H\dot{\sigma}-\frac{\partial V\left(\phi,\sigma\right)}{\partial\sigma}=0.$ (25)
In quintom cosmology, the effective dark energy is attributed to the canonical and phantom fields, and thus its energy density and pressure are given by
$\rho_{DE}\equiv\rho_{\phi}+\rho_{\sigma}=\frac{1}{2N^{2}}\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)+V\left(\phi,\sigma\right)$ (26)
$p_{DE}\equiv p_{\phi}+p_{\sigma}=\frac{1}{2N^{2}}\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)-V\left(\phi,\sigma\right),$ (27)
while its equation-of-state parameter is
$w_{DE}\equiv\frac{p_{DE}}{\rho_{DE}}=\frac{\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)-2N^{2}V\left(\phi,\sigma\right)}{\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)+2N^{2}V\left(\phi,\sigma\right)}.$ (28)
IV Dynamical equivalence of $f(Q)$ cosmology with Quintom scenario
In this section we will show the dynamical equivalence of $f(Q)$ cosmology, in the minisuperspace framework, with the quintom scenario. Considering that in (19) the lapse function is $N=\bar{N}\varphi$, it follows
$L\left(\Gamma^{B}\right)=-\frac{3}{\bar{N}}a\dot{a}^{2}-\frac{3}{2\bar{N}}a^{3}\frac{\dot{\varphi}}{\varphi}\dot{\psi}-\bar{N}a^{3}\left(\varphi V\left(\varphi\right)\right).$ (29)
Defining $\Phi=\ln\varphi$ we acquire
$L\left(\Gamma^{B}\right)=-\frac{3}{\bar{N}}a\dot{a}^{2}-\frac{3}{2\bar{N}}a^{3}\dot{\Phi}\dot{\psi}-\bar{N}a^{3}\left(e^{\Phi}V\left(\Phi\right)\right).$ (30)
Finally, introducing
$\Phi$ $=$ $\frac{1}{\sqrt{3}}\left(\phi+\sigma\right)$ (31)
$\psi$ $=$ $\frac{1}{\sqrt{3}}\left(\sigma-\phi\right),$ (32)
equivalently we have
$L\left(\Gamma^{B}\right)=-\frac{3}{\bar{N}}a\dot{a}^{2}+\frac{1}{2\bar{N}}a^{3}\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)-\bar{N}a^{3}\hat{V}\left(\phi+\sigma\right),$ (33)
where $\hat{V}\left(\phi+\sigma\right)=e^{\left(\phi+\sigma\right)}V\left(\phi+\sigma\right)$. In this case, the first Friedmann equation is written as
$3\bar{H}^{2}=\frac{1}{2\bar{N}^{2}}\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)+\hat{V}\left(\phi+\sigma\right),$ (34)
with $\bar{H}=\frac{1}{\bar{N}}\frac{\dot{a}}{a}$, and where we have defined $H=\frac{1}{\phi+\sigma}\bar{H}$. Consequently, we can define the effective dark energy sector with energy density and pressure
$\!\!\!\!\!\!\!\!\!\!\!\!\rho_{DE}^{Q}\equiv\rho_{\phi}+\rho_{\sigma}=\frac{1}{2\bar{N}^{2}}\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)+\hat{V}\left(\phi+\sigma\right)$ (35)
$\!\!\!\!\!\!\!\!\!\!\!\!p_{DE}^{Q}\equiv p_{\phi}+p_{\sigma}=\frac{1}{2\bar{N}^{2}}\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)-\hat{V}\left(\phi+\sigma\right),$ (36)
such that the equation-of-state parameter is defined as
$w_{DE}\equiv\frac{p_{DE}}{\rho_{DE}}=\frac{\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)-2N^{2}V\left(\phi,\sigma\right)}{\left(\dot{\phi}^{2}-\dot{\sigma}^{2}\right)+2N^{2}V\left(\phi,\sigma\right)}.$ (37)
As we see, $f(Q)$ cosmology with the second connection gives rise to a two-field model. In particular, one field comes from the non-linear $f(Q)$ form, and the other one arises from the extra connection field $\gamma$, while their combination yields the final fields, $\phi$ and $\sigma$, that appear in the Friedmann equation. Interestingly enough, we see that one field appears in the standard canonical way, while the other appears with opposite kinetic energy, ans thus exhibiting effectively a phantom behavior. Thus, $f(Q)$ cosmology with the second connection is dynamically equivalent with a quintom-like model. We use the term quintom-like and not quintom, since we have obtained a nonminimal coupling through $H=\frac{1}{\phi+\sigma}\bar{H}$, i.e. we obtain an effective gravitational constant proportional to $1/({\phi+\sigma})^{2}$. However, dynamically any trajectory of the basic quintom model is equivalent to the $f\left(Q\right)$ cosmology. The observable physics may be different, except in the asymptotic limit where $\phi+\sigma$ is constant, that is $Q$ becomes constant and the limit of General Relativity is recovered.
We mention here that one can obtain a full equivalence between $f(Q)$ cosmology and the quintom scenario, if he extends to nonminimal coupled quintom. In this case, using terms of the form $f(\phi)R$ and $f(\sigma)R$ gives rise to Friedmann equations with effective gravitational constant, where $H^{2}$ is multiplied by functions of $\phi$ and $\sigma$ Setare:2008pc . Additionally, one can straightforwardly consider the matter sector too, without or with nonminimal couplings, or interactions with the scalar fields.
Nevertheless, the important result that we desire to point out in this first paper on the subject, is that we have provided a way to obtain a phantom field from first principles, namely from a non-trivial connection. Indeed, in the rich literature of phantom and quintom fields and their applications, the phantom field is introduced in the action ad hoc, and thus one typically faces all the known problems that the phantom fields could have at the quantum level, since the energy is unbounded from below. However, in this work we showed that what appears effectively as a phantom field, is the combination of the well-behaved $f(Q)$ scalar field together with the well-behaved connection field $\gamma$ of the non-trivial connection. In summary, this may open a new way to handle the phantom fields, namely as effective realizations of the connection-related canonical fields.
Regarding the dynamical degrees of freedom of the theory, we mention that in the connections II and III the nonmetricity scalar has a boundary term contribution which depends on $\gamma$. Due to this boundary term, the nonmetricity has second-order derivative terms of the scale factor and of the derivative of $\gamma$. Consequently, a nonlinear function $f\left(Q\right)$ leads to an analogy of a fourth-order theory, with the additional scalar $\dot{\psi}=\gamma$. Hence, by using a Lagrange multiplier we were able to re-write the fourth-order theory as a second-order one, by increasing the dependent variables (introducing a second scalar field), and due to the structure of the theory the combination of the two fields appears effectively as a phantom field. Nevertheless, for the connection I, the non-metricity does not have any boundary term contribution, and it depends only on first-order derivatives of the scale factor, leading finally to a second-order theory of gravity.
V Conclusion
We showed that $f(Q)$ cosmology is dynamical equivalent with a quintom-like model. $f(Q)$ gravity is a modified theory of gravity that uses the non-metricity tensor as the fundamental geometric quantity which describes the gravitational field, and its cosmological applications are very rich and have been widely investigated. On the other hand, the quintom scenario is based on the simultaneous consideration of one canonical and one phantom scalar field, and its cosmological phenomenology is very interesting since it can naturally realize the phantom-divide crossing for the dark energy sector. Nevertheless, the introduction of the phantom field in phantom and quintom cosmology is ad hoc, and the fact that the energy of the phantom field is unbounded from below makes it potentially unstable at the quantum level.
We formulated $f(Q)$ cosmology in the minisuperspace approach, considering also the non-trivial connection, namely Connection II of the literature. In this case the scalar field arising from the non-linear $f(Q)$ form, and the scalar field associated to the non-trivial connection, are combined to give one canonical and one phantom field in the minisuperspace Lagrangian. In other words, the combination of two well-behaved fields appears effectively as a phantom field. The resulting Lagrangian and Friedmann equations are dynamically equivalent with a quintom-like scenario.
In summary, although in the standard approach to phantom and quintom scenario, the phantom field is added by hand, which then may raise issues concerning the stability, in the present analysis we showed that the phantom fields can arise as effective realizations of the connection-related canonical fields, and thus the quintom scenario is just a particular case of $f(Q)$ cosmology. This correspondence may open a new avenue to handle phantom fields and their applications in a cosmological framework.
We close this work with the following comments. $f\left(Q\right)$ gravity is known to exhibit some pathologies, investigated in detail in Gomes:2023tur , where it was found that strong coupling and ghosts appear. However, in Guzman:2024cwa it was found that connection II offers possibilities for a realistic cosmological description, assuming a specific set of initial conditions. The scalar-field formalism, that we presented in this work, can be used to investigate further this subject. Additionally, it is worth to investigate the conformal transformation of the theory. Finally, one powerful approach that allows to study the existence of pathologies in a theory is the projective symmetry BeltranJimenez:2019acz ; BeltranJimenez:2020sqf , since projective invariance is a sufficient condition for the absence of ghost-like instabilities. Hence, examining the projective symmetry of $f(Q)$ theory with connection II is both interesting and necessary, and can reveal important information about the structure of the theory. Since such a full analysis is out of the scopes of this work, it will be performed in a future project.